Suppose the average price of gasoline for a city in the United States follows a continuous uniform distribution with a lower bound of $3.50 per gallon and an upper bound of $3.80 per gallon. What is the probability a randomly chosen gas station charges between $3.70 and $3.90 per gallon?

A. 0.3000

B. 1.0000

C. 0.6667

D. 0.3333
Answer #1

Solution :

Given that,

a = 3.50

b = 3.80

P(c < x < d) = (d - c) / (b - a)

P(3.70 < x < 3.90) = (3.90 - 3.70) / (3.80 - 3.50) = 0.2 / 0.3 = 0.6667

Probability = 0.667

Option C)